# Credit Card Fraud Detection Using Advanced Transformer Model

### v2

@@ I Introduction
A major technical challenge lies in the fact that fraudulent transactions account for only a small portion of credit card payments, making it challenging for machine learning algorithms to detect such sensitive behavior[29]. As the most powerful and capable model to date, the Transformer large model exhibits revolutionary identification capabilities, making it particularly suitable for recognizing fraudulent data in small samples.
@@
In recent years, various machine learning techniques have been applied to tackle the problem of fraud detection. Neural Networks (NN)[7], for instance, have been widely used due to their ability to learn complex patterns and relationships from large datasets [22]. NNs can automatically extract features from raw data and create a hierarchical representation, enabling them to detect fraudulent activities with high accuracy [1].
@@
Another popular approach is XGBoost, an ensemble learning method that combines multiple decision trees to create a robust and accurate model [5]. XGBoost has proven to be effective in fraud detection tasks, as it can handle imbalanced datasets and capture non-linear relationships between features[36].
@@
TabNet, a more recent addition to the fraud detection toolkit, is a deep learning architecture specifically designed for tabular data [3]. It utilizes an attentive mechanism to learn the importance of each feature and creates interpretable explanations for its predictions. TabNet has shown promising results in fraud detection, particularly in scenarios where interpretability is crucial[14].
@@
Despite the success of these techniques, they often face challenges when dealing with complex, high-dimensional data and long-range dependencies. This is where Transformer models, originally developed for natural language processing tasks, have shown immense potential[6]. Transformers can effectively capture intricate patterns and relationships within the data, thanks to their self-attention mechanism. By adapting Transformers for fraud detection, we can leverage their ability to handle complex data structures and create more accurate and robust models.
@@
Moreover, Transformers can be pre-trained on large, unlabeled datasets and then fine-tuned for specific fraud detection tasks. This transfer learning approach allows the models to learn general representations of the data and reduces the need for extensive labeled datasets. Transformers have the potential to outperform traditional methods and provide a more efficient and effective solution for fraud detection in various domains. Therefore, in this study, we aim to thoroughly explore the feasibility of applying Transformers to fraud detection.
@@
The remainder of this paper is organized as follows. In Section II, we provide a comprehensive review of the related research that has informed the objectives and methodological choices of the current study. This review covers key contributions in areas such as statistical modeling, neural networks, and machine learning approaches to credit card fraud detection[38].
@@ II Related Work
Similar work has been done in other deep learning biological areas such as [17] fine-tuning segment anything to obtain accurate forged regions [18] proposing a new style sampling module that improves the detection accuracy and generalization of the anomalous face classification task
In the realm of model selection, the success of large language models in other domains, as showcased by the research of Wang et al[33]., has provided important guidance. To ensure the efficacy of training and testing strategies, the data cleaning and processing methods developed by Wei et al[34]. have been invaluable. Shen et al[27]s work on efficient classification models has offered valuable experience in model selection, while Ding et al.’s[9] approach to data filtering, segmentation, and augmentation has provided useful insights for addressing sparse data challenges.
@@ III Methodology
In Section III, we provide a detailed account of the methodology employed in this work. This includes a thorough discussion of the data processing techniques utilized, such as ensuring an equally distributed dataset. We also examine the relevant prior research on feature engineering and dimensionality reduction, and analyze the impact these techniques had on the data used in our study. By comprehensively documenting these methodological components, we aim to establish transparency and enable the replication of our research approach.
The primary focus of our methodology lies in accurately identifying instances of credit card fraud in datasets where fraudulent samples are exceedingly rare. We conducted comprehensive evaluations and explorations of various model architectures, accompanied by detailed and thorough data analysis and preprocessing.[35] Additionally, we selected relevant loss functions and evaluation methods to ensure the robustness and effectiveness of our approach.
@@ III-A Dataset Introduction
The dataset used in this study consists of European credit card transaction data. To ensure the timeliness of the experiments and adapt to the latest payment environment, we utilized the most recent data available, encompassing over 550,000 transaction records of European credit card users up to 2023. These data contain the latest processed variables, ranging from V1 to V28. To enhance the informative value of the data and increase the scientific validity of the comparisons, we additionally incorporated European payment data from 2013, thereby increasing the sample size and improving the meaningfulness of our experimental comparisons. Among the dataset’s 284,804 transactions, 492 were identified as fraudulent, accounting for 0.172% of the total. All these data originate from the same source and possess variables V1 to V28, which have been processed and converted to float values.
@@ III-B1 Equally Distributing and Correlating
To generate an equally distributed dataset, the initial step involves displaying textual output showcasing the proportions of class distribution to confirm the balance of the dataset[20]. Following this, a bar plot is generated utilizing Seaborn’s countplot function to visually depict the distribution of classes. This visualization illustrates an equal distribution of classes, thereby underscoring the effectiveness of the subsampling technique in achieving class balance[21].
@@ III-B2 Feature Correlation Analysi
In our analysis, we meticulously compared the original imbalanced dataset with its subsampled counterpart, unearthing a plethora of advantages stemming from data balancing[40]. This encompassed a noticeable boost in model performance, mitigated susceptibility to overfitting, heightened predictive stability, and augmented model interpretability. The correlation efficacy visibly escalated as evidenced by the comparison, underscoring the tangible benefits derived from this approach. Correspond info please check Fig1 and Fig2.
Figure 1: Figure 4: Imbalanced Correlation matrix
Figure 2: Figure 5: Subsample correlation matrix
@@ III-B3 Outlier Detection
Calculate quartiles and interquartile range: First, calculate the 25th percentile (Q1) and 75th percentile (Q3) of the feature values. Then, calculate the interquartile range (IQR = Q3 - Q1). Determine the upper and lower bounds for outliers: Use 1.5 times the IQR to determine the upper and lower bounds for outliers, i.e., the lower bound is Q1 - 1.5 * IQR, and the upper bound is Q3 + 1.5 * IQR. Remove outliers: Consider feature values that are lower than the lower bound or higher than the upper bound as outliers and remove these outliers from the dataset. For example, for feature V14, we first calculate its quartiles and interquartile range, then determine the upper and lower bounds for outliers, and finally remove outliers that exceed these ranges. The same method is also applied to features V12 and V10. To verify the effect of outlier handling, we again conducted a visual analysis of the processed data, using box plots to show the distribution of features V14, V12, and V10 in fraudulent and non-fraudulent transactions. Box plots not only display the central tendency and dispersion of the data but also visually show outliers as Fig 4.
Figure 4: n
@@ III-B4 Dimensionality Reduction
T-SNE (t-Distributed Stochastic Neighbor Embedding) is a non-linear dimensionality reduction technique particularly well-suited for embedding high-dimensional data into a space of two or three dimensions. It models each high-dimensional object by a two- or three-dimensional point in such a way that similar objects are modeled by nearby points and dissimilar objects are modeled by distant points with high probability[31]. T-SNE excels in preserving the nonlinear structures and local relationships within high-dimensional data. Consequently, T-SNE offers superior advantages in data visualization and clustering analysis.
@@
PCA (Principal Component Analysis) is a linear dimensionality reduction technique that transforms the original data into a new coordinate system where the greatest variances by any projection of the data come to lie on the first coordinates (called principal components), the second greatest variances on the second coordinates, and so on. PCA is efficient and widely used for reducing the dimensionality of datasets while preserving as much variability as possible[2].
@@
Truncated SVD (Singular Value Decomposition) is a matrix factorization technique that decomposes a matrix into three other matrices. It is particularly useful for sparse data and is often used in the context of Latent Semantic Analysis (LSA) in natural language processing. Unlike PCA, Truncated SVD does not center the data before performing the decomposition, making it suitable for term-document matrices and other non-centered data[13].
@@
Figure 5: Comparison of T-SNE, PCA, and Truncated SVD
@@ III-C1 Transformer Structure
The Self-Attention Mechanism is utilized to compute attention weights for each position in the input sequence, capturing dependencies between different positions. In the Transformer Encoder[41], the Self-Attention Mechanism is referred to as Multi-Head Attention. Below is the mathematical formulation of Multi-Head Attention:
@@
Next, we compute the attention weights ($A$) by taking the dot product of Query and Key, followed by scaling and softmax normalization:
@@
Finally, we multiply the attention weights with the values and then perform a weighted sum to obtain the output $Z$ of the Self-Attention Mechanism[23]:
$Z=AV$
The Multi-Head Attention mechanism calculates the outputs of multiple attention heads by using independent Query, Key, and Value projection matrices. These outputs are concatenated along the last dimension to obtain the final output of the Self-Attention Mechanism.
@@
In the self-attention mechanism, we use the input sequence as queries, keys, and values, and compute attention weights for each position with respect to other positions[24]. Then, these attention weights are used to weight-sum the values of the input sequence, resulting in the output of the self-attention mechanism[37].
@@
$\text{EncoderLayer}(X)$ $=\text{LayerNorm}(X+\text{Attention}(X,X,X))$
@@
Here is the structure of our implemented model Figure 6:
Figure 7: Enter Caption
@@ IV-A Evaluation Metrics
In this experiment, our evaluation metrics include Precision, Recall, F1-score, and ROC AUC. These metrics provide different perspectives when assessing the classification results of the model.
@@
Precision refers to the proportion of true positive samples among all[11] samples predicted as positive by the model. It can be represented by the following formula:
$\text{Precision}=\frac{TP}{TP+FP}$
where $TP$ (True Positives) is the number of true positive samples, and $FP$ (False Positives) is the number of false positive samples.
@@
Recall indicates the proportion of true positive samples among[25] all actual positive samples. It can be represented by the following formula:
$\text{Recall}=\frac{TP}{TP+FN}$
Here, $FN$ (False Negatives) represents the number of false negative samples.
@@
The F1-score is a metric that strikes a balance between precision and recall by calculating their harmonic mean. It is computed using the following formula:
[…]
@@
The Receiver Operating Characteristic Area Under the Curve is employed as a performance metric for classification tasks across various threshold settings. The ROC curve is plotted by placing the true positive rate (Recall) on the y-axis and the false positive rate (FPR) on the x-axis. The AUC[30], or Area Under the Curve, indicates how well the model distinguishes between classes, with higher values signifying superior performance. Essentially, it measures the entire two-dimensional area underneath the ROC curve.
Interpreting the ROC AUC involves understanding it as the likelihood that a randomly chosen positive instance[12] will be ranked higher by the classifier than a randomly chosen negative instance. This metric provides a thorough evaluation by considering both sensitivity (Recall) and specificity.
The selection of these metrics is contingent upon the specific problem and its requirements. For instance, scenarios that prioritize the reduction of false negatives will emphasize recall. Conversely, other situations might value precision more highly. Both the F1-score and ROC AUC deliver extensive evaluations that account for precision, recall, and overall model efficacy.
@@ IV-B Results
To fairly and accurately select the best fraud detection model, we carefully curated several different classification models. The models designed for this experiment include Logistic Regression, K-Nearest Neighbors, Support Vector Machine, Decision Tree Classifier, and Neural Network. Additionally, to ensure that the models used in our experiments remain highly practical in comparison to the currently used models in the same scenarios, we included comparisons with the well-known and excellent model XGBoost and the deep learning-based TabNet.
@@
TABLE I: Experiment with 2023 Data
@@
Figure 8: Fig 8 ROC AUC
@@
Figure 9: ROC AUC Scores of Various Classifiers on 2013 Data
@@
Secondly, the Transformer model’s F1-score on the 2013 data also reaches 0.998, significantly surpassing other models. This further confirms the Transformer’s superiority in balancing Precision and Recall [8]. It is worth noting that TabNet performs poorly on the 2013 data, with an F1-score of only 0.67. This may be attributed to TabNet’s sensitivity to changes in data distribution, resulting in unstable performance across different time periods [3].
Moreover, the Transformer model achieves an ROC AUC of 0.98 on the 2013 data, which is only slightly lower than XGBoost’s 0.96. However, considering its performance in other metrics, the Transformer still demonstrates the best overall performance. This further validates the Transformer’s advantage in capturing complex feature interactions and long-range dependencies, enabling it to maintain consistently high performance on data from different time periods [10].
@@ V Conclusion
The success of this experiment can be attributed to our extensive efforts in understanding the underlying principles of the Transformer and adjusting its Encoder layers. Despite the considerable challenge of adapting the Transformer[33], initially designed for natural language processing tasks, into a classification model suitable for fraud detection, our model ultimately lived up to expectations, demonstrating outstanding accuracy across all evaluation metrics. This result not only proves the superior performance of the Transformer model in the fraud detection domain but also provides strong support for its application in other fields.
As we move forward, our focus will remain on investigating the potential of Transformer models in fraud detection and beyond. Our goal is to create models that are not only more efficient and accurate but also more resilient to various challenges. Moreover, we believe that the findings of this experiment will serve as a valuable resource for professionals in the industry, encouraging the adoption and advancement of AI-driven solutions in the realm of financial security and other critical areas.

### v3

@@ I Introduction
A major technical challenge lies in the fact that fraudulent transactions account for only a small portion of credit card payments, making it challenging for machine learning algorithms to detect such sensitive behavior[42, 57]. As the most powerful and capable model to date, the Transformer large model exhibits revolutionary identification capabilities, making it particularly suitable for recognizing fraudulent data in small samples.
@@
In recent years, various machine learning techniques have been applied to tackle the problem of fraud detection. Neural Networks (NN)[8], for instance, have been widely used due to their ability to learn complex features and relationships from large datasets [26]. NNs can automatically extract features from raw data and create a hierarchical representation, enabling them to detect fraudulent activities with high accuracy [1].
@@
Another popular approach is XGBoost, an ensemble learning method that combines multiple decision trees to create a robust and accurate model [6, 3]. XGBoost has proven to be effective in fraud detection tasks, as it can handle imbalanced datasets and capture non-linear relationships between features[51].
@@
TabNet, a more recent addition to the fraud detection toolkit, is a deep learning architecture specifically designed for tabular data [4]. It utilizes an attentive mechanism to learn the importance of each feature and creates interpretable explanations for its predictions. TabNet has shown promising results in fraud detection, particularly in scenarios where interpretability is crucial[16, 27].
@@
Despite the success of these techniques, they often face challenges when dealing with complex, high-dimensional data and long-range dependencies. This is where Transformer models, originally developed for natural language processing tasks, have shown immense potential[7]. Haowei’s work[37] has demonstrated the superiority of transformer-based models over traditional statistical approaches in accurately predicting heart rate time series, effectively capturing complex temporal dependencies and non-linear relationships inherent in physiological data. Despite the success of these techniques, they often face challenges when dealing with complex, high-dimensional data and long-range dependencies. This is where Transformer models, originally developed for natural language processing tasks, have shown immense potential[7, 58]. Transformers can effectively capture intricate patterns and relationships within the data, thanks to their self-attention mechanism. By adapting Transformers for fraud detection, we can leverage their ability to handle complex data structures and create more accurate and robust models.
@@
Moreover, Transformers can be pre-trained on large, unlabeled datasets and then fine-tuned for specific fraud detection tasks. This transfer learning approach allows the models to learn general representations of the data and reduces the need for extensive labeled datasets[35]. Transformers have the potential to outperform traditional methods and provide a more efficient and effective solution for fraud detection in various domains. Therefore, in this study, we aim to thoroughly explore the feasibility of applying Transformers to fraud detection.
@@
The remainder of this research is written as follows. In Section II, we provide a comprehensive review of the related research that has informed the objectives and methodological choices of the current study. This review covers key contributions in areas such as statistical modeling, NN, and machine learning approaches to credit card fraud detection[53].
@@ II Related Work
Similar work has been done in other deep learning biological areas such as [21, 19] fine-tuning segment anything to obtain accurate forged regions [22] proposing a new style sampling module that improves the detection accuracy and generalization of the anomalous face classification task
In the realm of model selection, the success of large language models in other domains, as showcased by the research of Wang et al[47, 34]., has provided important guidance. To ensure the efficacy of training and testing strategies, the data cleaning and processing methods developed by Wei et al[48]. have been invaluable. Shen et al[39]s work on efficient classification models has offered valuable experience in model selection, while Ding et al.’s[10] approach to data filtering, segmentation, and augmentation has provided useful insights for addressing sparse data challenges.
In their groundbreaking paper, Ni et al. (2023) present a comprehensive comparison between traditional models and deep learning approaches for heart rate time series forecasting[37].Their research demonstrates that transformer-based models, particularly PatchTST, significantly outperform traditional models like ARIMA and Prophet in predicting heart rate dynamics[37].
@@ III Methodology
In Section III, we provide a detailed account of the methodology employed in this work. This includes a thorough discussion of the data processing techniques utilized, such as ensuring an equally distributed dataset. We also examine the relevant prior research on feature engineering and dimensionality reduction, and analyze the impact these techniques had on the data used in our study[59, 56]. By comprehensively documenting these methodological components, we aim to establish transparency and enable the replication of our research approach.
The primary focus of our methodology lies in accurately identifying instances of credit card fraud in datasets where fraudulent samples are exceedingly rare. We conducted comprehensive evaluations and explorations of various model architectures[60], accompanied by detailed and thorough data analysis and preprocessing.[50, 18] Additionally, we selected relevant loss functions and evaluation methods to ensure the robustness and effectiveness of our approach. The study[37] has significantly influenced our research, guiding the development of advanced predictive models in our work to handle complex temporal dependencies and non-linear relationships.
@@ III-A Dataset Introduction
The dataset used in this study consists of European credit card transaction data. To ensure the timeliness of the experiments and adapt to the latest payment environment, we utilized the most recent data available, encompassing over 550,000 transaction records of European credit card users up to 2023. These data contain the latest processed variables, ranging from V1 to V28. To enhance the informative value of the data and increase the scientific validity of the comparisons, we additionally incorporated European payment data from 2013, thereby increasing the datasource size and improving the meaningfulness of our experimental comparisons. Among the dataset’s 284,804 transactions, 492 were identified as fraudulent, accounting for 0.172% of the total. All these data originate from the same source and possess variables V1 to V28, which have been processed and converted to float values.
@@ III-B1 Equally Distributing and Correlating
To generate an equally distributed dataset, the initial step involves displaying textual output showcasing the proportions of class distribution to confirm the balance of the dataset[49]. Following this, a bar plot is generated utilizing Seaborn’s countplot function to visually depict the distribution of classes. This visualization illustrates an equal distribution of classes, thereby underscoring the effectiveness of the subsampling technique in achieving class balance.
@@ III-B2 Feature Correlation Analysi
In our analysis, we meticulously compared the original imbalanced dataset with its subsampled counterpart, unearthing a plethora of advantages stemming from data balancing[55, 61]. This encompassed a noticeable boost in model performance, mitigated susceptibility to overfitting, heightened predictive stability, and augmented model interpretability. The correlation efficacy visibly escalated as evidenced by the comparison, underscoring the tangible benefits derived from this approach. Correspond info please check Fig1 and Fig2.
Figure 1: Imbalanced Correlation matrix.
Figure 2: Sub-sampled correlation matrix.
@@ III-B3 Outlier Detection
Calculate quartiles and interquartile range: First, calculate the 25th percentile (Q1) and 75th percentile (Q3) of the feature values. Then, calculate the interquartile range (IQR = Q3 - Q1). Determine the upper and lower bounds for outliers: Use 1.5 times the IQR to determine the upper and lower bounds for outliers, i.e., the lower bound is Q1 - 1.5 * IQR, and the upper bound is Q3 + 1.5 * IQR. Remove outliers: Consider feature values that are lower than the lower bound or higher than the upper bound[14] as outliers and remove these outliers from the dataset. For example, for feature V14, we first calculate its quartiles and interquartile range, then determine the upper and lower bounds for outliers, and finally remove outliers that exceed these ranges. The same method is also applied to features V12 and V10. To verify the effect of outlier handling, we again conducted a visual analysis of the processed data, using box plots to show the distribution of features V14, V12, and V10 in fraudulent and non-fraudulent transactions. Box plots not only display the central tendency and dispersion of the data but also visually show outliers as Fig 4.
Figure 4: Outlier Removal for V14 and V12 feature.
@@ III-B4 Dimensionality Reduction
T-Distributed Stochastic Neighbor Embedding, is a technique for reducing the dimensionality of high-dimensional data, particularly when the goal is to represent the data in a low-dimensional space of two or three dimensions. Unlike linear dimensionality reduction methods, T-SNE is capable of capturing and preserving the complex nonlinear structures and local relationships inherent in high-dimensional data[45]. The algorithm works by representing each high-dimensional data point as a corresponding low-dimensional point in a way that ensures similar objects are represented by nearby points and dissimilar objects are represented by distant[36] points with a high degree of probability. This unique ability to maintain the intricate patterns and local similarities of the original data makes T-SNE an invaluable tool for data visualization and clustering analysis, providing researchers with a clear and intuitive understanding of the underlying structure and relationships within the high-dimensional data.
@@
PCA (Principal Component Analysis) is a linear dimensionality reduction technique that[36] transforms the original data into a new coordinate system where the greatest variances by any projection of the data come to lie on the first coordinates (called principal components), the second greatest variances on the second coordinates, and so on. PCA is efficient and widely used for reducing the dimensionality of datasets while preserving as much variability as possible[2].
@@
Truncated SVD (Singular Value Decomposition) is technique that decomposes a matrix into three other matrices. It is particularly useful for sparse data and is often used in the context of Latent Semantic Analysis (LSA) in natural language processing. Unlike PCA, Truncated SVD does not center the data before performing the decomposition, making it suitable for term-document matrices and other non-centered data[15].
@@
Figure 5: Comparison of T-SNE, PCA, and Truncated SVD.
@@ III-C1 Transformer Structure
The Self-Attention Mechanism is utilized to compute attention weights for each position in the input sequence, capturing dependencies between different positions. In the Transformer Encoder, the Self-Attention Mechanism is referred to as Multi-Head Attention. Below is the mathematical formulation of Multi-Head Attention:
@@
Next, we compute the attention weights ($A$) by taking the dot product of Query and Key[41], followed by scaling and softmax normalization:
@@
Finally, we multiply the attention weights with the values and then perform a weighted sum to obtain the output of the Self-Attention Mechanism[29]:
$Attention(X)=AV$
The Multi-Head Attention mechanism calculates the outputs of multiple attention heads by using independent Query, Key, and Value projection matrices. These outputs are concatenated along the last dimension to generate the final result of the Self-Attention Mechanism.
@@
In the self-attention mechanism, we use the input sequence as queries, keys, and values and compute attention weights for each position with respect to other positions[30]. Then, these attention weights are used to weight-sum the values of the input sequence, resulting in the output of the self-attention mechanism[52].
@@
$\text{EncoderLayer}(X)$ $=\text{LayerNorm}(X+\text{Attention}(X))$
@@
Here is the Structure of our implemented model. Figure 6:
Figure 7: Structure of attention module.
@@ IV-A Evaluation Metrics
We use Precision, Recall, F1-score, and ROC AUC to assess our classification model. These metrics offer a thorough evaluation from different perspectives.
@@
Precision calculates the rate of true positive predictions among all positive predictions, indicating the model’s accuracy in identifying positives[12].
@@
Recall, or sensitivity, calculates the proportion of true positives detected among all actual positives, showing the model’s ability to capture positive cases[32].
@@
The F1-score provides a balanced measure of the model’s accuracy in predicting positives and capturing actual positives.
@@
ROC AUC evaluates the model’s ability to distinguish between classes at various thresholds. It represents the area under the ROC curve, where a higher AUC indicates better discrimination between positive and negative instances[43][13].
The choice of metrics depends on the problem’s requirements, with some scenarios prioritizing recall and others precision. The F1-score and ROC AUC offer balanced evaluations, considering both aspects.
@@ IV-B Results
To fairly and accurately select the best fraud detection model, we carefully curated several different classification models. The models designed for this experiment include Logistic Regression, K-Nearest Neighbors, SVC, Decision Tree Classifier, and Neural Network. Additionally, to ensure that the models used in our experiments remain highly practical in comparison to the currently used models in the same scenarios, we included comparisons with the well-known and excellent model XGBoost and the deep learning-based TabNet.
@@
TABLE I: Experiment with 2023 Data.
@@
Figure 8: ROC and AUC Scores of Various Classifiers on 2023 Data.
@@
Figure 9: ROC AUC Scores of Various Classifiers on 2013 Data.
@@
Secondly, the Transformer model’s F1-score on the 2013 data also reaches 0.998, significantly surpassing other models. This further confirms the Transformer’s superiority in balancing Precision and Recall [9, 44]. It is worth noting that TabNet performs poorly on the 2013 data, with an F1-score of only 0.67. This may be attributed to TabNet’s sensitivity to changes in data distribution, resulting in unstable performance across different time periods [4].
Moreover, the Transformer model achieves an ROC AUC of 0.98 on the 2013 data, which is only slightly lower than XGBoost’s 0.96. However, considering its performance in other metrics, the Transformer still demonstrates the best overall performance. This further validates the Transformer’s advantage in capturing complex feature interactions and long-range dependencies, enabling it to maintain consistently high performance on data from different time periods [11, 25].
@@ V Conclusion
The success of our experiment can be attributed to our extensive efforts in understanding the underlying principles of the Transformer and adjusting its Encoder layers. Despite the considerable challenge of adapting the Transformer[47, 37], initially designed for natural language processing tasks, into a classification model suitable for fraud detection, our model ultimately lived up to expectations, demonstrating outstanding accuracy across all evaluation metrics. This result not only proves the superior performance of the Transformer model in the fraud detection domain but also provides strong support for its application in other fields[31, 38].
In the future research, our focus will remain on investigating the potential of Transformer models in fraud detection and beyond[28, 24]. Our goal is to create models that are not only more efficient and accurate but also more resilient to various challenges. Moreover, we believe that the findings of this experiment will serve as a valuable resource for professionals in the industry, encouraging the adoption and advancement of AI-driven solutions in the realm of financial security and other critical areas.